OpenStudy (anonymous):

Several volcanoes have been observed erupting on the surface of a planet's closest moon. Suppose that material ejected from one of these volcanoes reaches a height of 4.20 km after being projected straight upward with an initial speed of 205 m/s. If the radius of this moon is 4211 km, outline a strategy that allows you to calculate the mass of Io.

OpenStudy (anonymous):

How can I calculate the mass of the planet, if I don't have a mass of a second body..

OpenStudy (experimentx):

use distance formulas .. to calculate acceleration ... use acceleration due to gravity to calculate mass of moon

OpenStudy (anonymous):

I'm having trouble finding which distance formula to use.

OpenStudy (experimentx):

v^2 = u^2+2as V = 0, s = 4200

OpenStudy (anonymous):

Is it d = vt + (1/2)at^2?

OpenStudy (experimentx):

don't bother with time

OpenStudy (anonymous):

In your equation, what is u, and what is s? And how did you get 4200?

OpenStudy (experimentx):

u is given ... initial velocitu ... s is distance in meters ... change km in m

OpenStudy (anonymous):

or sorry, u is 205m/s?

OpenStudy (jamesj):

Wait a second. Using Newton's Universal Law of Gravitation, what is the expression for the acceleration due to gravity, an expression that involves the moon's mass?

OpenStudy (anonymous):

All I know is g=9.8, but I don't think that applies here..

OpenStudy (jamesj):

It most emphatically does not. You will need Newton's ULG.

OpenStudy (experimentx):

no ... that does not apply here

OpenStudy (experimentx):

it has a new value ... find it using above formula

OpenStudy (anonymous):

So you mean F=Gm1m2/r^2? I don't know how to use that here, because I don't have a second mass.

OpenStudy (experimentx):

and u is 205

OpenStudy (anonymous):

but how do I find v?

OpenStudy (experimentx):

change F = m1 a ... and cancel out m1 ... you will have m2 ,,, which is to be found out.

OpenStudy (anonymous):

hmm

OpenStudy (experimentx):